OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x^2+10x+14=0

OpenStudy (anonymous):

x^2 + 10x + 14 = 0 x^2 + 10x = -14 x^2 + 10x + 25 = -14 + 25 x^2 + 10x + 25 = 11 (x + 5)^2 = 11 x + 5 = ±√11 x = ±√11 - 5